1. Use the method of reduction of order to find a second solution of the differential equation
t^2 y double prime + 10t y prime - 10y = 0, t > 0 given the solution y1 = t

A. y2 = t^-1
B. y2 = 1
C.y2 = t
D.y2 = t^-10
E.y2 = t^10
F y2 = t^11
G. y2 = t^-11
1
Expert's answer
2020-07-08T18:40:46-0400

Assume "y_2(t)=v(t)y_1(t)," then


"y_2'(t)=v'(t)y_1(t)+v(t)y_1'(t)=v'(t)t+v(t)"

"y_2''(t)=v''(t)t+v'(t)+v'(t)=v''(t)t+2v'(t)"

Plugging these into the differential equation gives


"t^2 (v''t+2v')+10t(v't+v)-10vt=0"

"t^3v''+2t^2v'+10t^2v'+10tv-10vt=0"

Since "t>0"

"tv''+12v'=0"

Let "v'=u," we have


"tu'+12u=0"

By separation of variables


"{u'\over u }=-{12\over t}"

"u=C_1t^{-12}"

Thus


"v'=C_1t^{-12}"

"v=-{1\over 11}C_1t^{-11}+C_2"

"v=C_3t^{-11}+C_2"

"y_2=C_3t^{-10}+C_2t"

Now we take "C_3=1, C_2=0"


"y_2=t^{-10}"

"D.\ \ y_2=t^{-10}"